# MotionFix: Text-Driven 3D Human Motion Editing

### v1

@@ Abstract
The focus of this paper is 3D motion editing. Given a 3D human motion and a textual description of the desired modification, our goal is to generate an edited motion as described by the text. The challenges include the lack of training data and the design of a model that faithfully edits the source motion. In this paper, we address both these challenges. We build a methodology to semi-automatically collect a dataset of triplets in the form of (i) a source motion, (ii) a target motion, and (iii) an edit text, and create the new MotionFix dataset. Having access to such data allows us to train a conditional diffusion model, TMED, that takes both the source motion and the edit text as input. We further build various baselines trained only on text-motion pairs datasets, and show superior performance of our model trained on triplets. We introduce new retrieval-based metrics for motion editing, and establish a new benchmark on the evaluation set of MotionFix. Our results are encouraging, paving the way for further research on finegrained motion generation. Code and models will be made publicly available.
@@
Figure 1: Motion editing: We introduce a semi-automatically collected dataset MotionFix that contains diverse types of editing such as modifying body parts, changing certain moments of the motion, or editing the speed or the style (top). Training on this dataset, we develop TMED, a text-driven motion diffusion model, enabling 3D human motion editing from natural language descriptions (bottom).
11footnotetext: Work was done while at MPI.
1 Introduction
@@
Human motion control is an essential component of the animation pipeline, and involves creating a motion, as well as editing it until the motion matches the desired outcome. Text descriptions have emerged as one of the prominent ways to control motion generation [36, 48, 17, 51]. However, due to the inherent ambiguity in high-level language instructions, the resulting generation may not necessarily correspond to the motion one has in mind. An animator may then need to edit the motion further. Motion editing is non-trivial, arguably more complex than static pose editing, and may involve multiple types of instructions, such as changing the speed of a motion, modifying the repetitions for cyclic actions, adjusting the posture of a particular body or modifying a certain temporal segment of a motion. In this work, we aim, given an initial source motion and an edit description, to generate a new motion that follows the source motion and edits it according to the text instruction.
@@
There are existing approaches that can modify body coordinates [23] and lower/upper limbs [51, 48]. However, they require manual selection of the body parts, which prevents making edits beyond local modifications such as the speed of the overall action. On the other hand, some methods have been proposed to add more fine-grained control to text-to-motion generation; examples include temporal [2, 52] and spatial [3, 54, 15] compositions. These methods can be repurposed for editing, but would be limited to addition or subtraction of actions. Consequently, all these works are limited to specific types of edits. Instead, our work considers unrestricted edits described by language instructions. To this end, we collect a manually annotated dataset, MotionFix, that supports training generative models for the task of text-driven motion editing.
@@
Constructing a dataset for 3D human motion editing is non-trivial. In contrast to text-based image editing, where methods such as InstructPix2Pix [5] exploit large text-to-image generation models [43] to automatically create training data, there exists no 3D motion generation model which generalizes faithfully to unrestricted text inputs. Besides, dynamic edits can be more complex than static image edits. In fact, PoseFix [10] is a successful example of a 3D human body pose editing dataset; however, the difference between static poses are mapped to text in a rule-based manner using joint distances, and this would not be applicable to dynamic motions. In this work, we take a different route and mine existing motion capture (MoCap) datasets to find suitable motion pairs automatically, for which the differences are then manually described by typing text. By not relying on a generative model, we ensure motion quality; and by annotating the text (which is relatively fast), we achieve unrestricted edits.
@@
The key challenge in our semi-automatic data curation pipeline is how to find motion pairs that are similar enough for a meaningful and concise edit text to describe the difference. The difference between the source and target motions should not be too large to avoid the annotator typing a too-complex text or describing entirely the target motion, discarding the source. The motions should have similarities for the annotator to potentially make reference to the source. Our solution is to employ the recent TMR motion embedding space [37] that effectively captures semantics, as well as sufficient details for the body dynamics, thanks to its contrastively and generatively trained motion encoder. To form our candidate pairs for annotation, for each motion in a large MoCap collection [32], we retrieve the top-ranked motions according to their embedding similarity. Using crowdsourcing, we collect textual annotations for these pairs. The resulting dataset, MotionFix, is the first text-based motion editing dataset, which contains different types of edits as can be seen in Figure 1 (top). Some edits involve a specific body part (e.g., the hand should “make the circular motion a bit wider”), others alter the overall body dynamics (e.g., “make the circle bigger and walk faster” when walking in a circle).
@@
To benchmark our model and compare against baselines, we introduce new metrics on the evaluation set of MotionFix. Following the commonly adopted retrieval-based metrics in text-to-motion generation benchmarks [17], we perform motion-to-motion retrieval and check how often the ground-truth target motion is in the top ranks. We also report the ranking of the source motion to evaluate the proximity to the source. While this metric should not be too high – otherwise there would be no edit – it gives an intuition on whether the generated motion deviates too much from the source. Our experiments demonstrate that our conditional model trained on triplets generates motions that are closer to the target, compared to strong baselines we build on top of state-of-the-art text-to-motion generation methods, which have only access to text-motion pairs for training.
@@ 2 Related Work
3D motion generation from text. In contrast to the relatively mature areas of text-to-image generation [43] and text-based image editing [5], language-based 3D human motion generation is at its infancy. Initial work employs VAEs [26] with action label conditioning using a small set of categories [16, 35]. With the introduction of recent text-motion datasets [41, 17, 30], there has been increased interest in conditioning the generation on free-form language inputs [36, 17, 18, 47, 54, 19]. Recently, diffusion models [20] have been successfully integrated [51, 48, 49, 44, 50], producing state-of-the-art results in text-to-motion generation.
@@
Several works focus on increased controllability in motion generation, going beyond a single textual input. Examples include enabling temporal compositionality (a series of motions) [2, 29, 44], spatial compositionality [3, 51] (simultaneous motions), and a unified framework of timeline control [38]. Diffusion-based models were shown to be suitable for fine-grained local control such as joint trajectories [23, 50, 44] or keyframes [7]. Our work is in similar spirit in terms of providing more control to users; however, in contrast to the above, our focus is motion editing using language.
@@
Language-based human body editing. There are numerous traditional methods for editing 3D humans to generate movies [6], imposing space or time constraints [8], interactively generating motions using procedural animation [28, 34] or physics-based approaches [40]. Recent work using language can be grouped into pose [24, 10] or motion [14, 15] editing. In FixMyPose [24], the focus is on editing athletic human poses in synthetic images. In [10], a text-based 3D human pose editing method is developed, enabled through the collection of the PoseFix dataset containing language descriptions of differences between pairs of poses. PoseFix builds on the previous work of PoseScript [9], where a dataset of pose descriptions are automatically collected through a rule-based approach. Unlike PoseFix that concentrates on static poses, our MotionFix dataset involves dynamic motions where the space of possible edits are much larger, necessitating a different approach to data collection.
@@
In terms of dynamic bodies, current motion editing approaches can be separated into three categories: (a) Style-based editing or motion style transfer exploits datasets that contain a small set of style labels such as ‘angry’ and ‘old’ [1, 33, 27]. This line of work, focuses mostly on copying the style of one motion onto another, typically performing the same action. (b) Part-based editing considers selecting a subset of the body. MDM [48] shows the potential of diffusion models to edit the upper/lower body by text-conditioned motion inpainting. Similarly, MotionDiffuse [51] and FLAME [25] manually specify body parts to edit them with text. More recently, CoMo [22] uses LLMs to produce edit texts and demonstrates promising results for part editing. (c) Among heuristic-based approaches [14, 15], AIFit [14] can edit domain-specific exercise poses from a pre-defined grammar and is focused on a limited set of cyclic motions from their Fit3D dataset. Iterative Motion Editing [15] relies on captioned source motions, which are passed through an LLM along with a pre-defined set of ‘Motion Editing Operators’ (MEOs), to detect which joints and frames should be edited. A pre-trained diffusion model is then used to infill these locations. In contrast with the prior work, we do not focus on a specific type of motion editing or any heuristics. Our MotionFix contains diverse edits as can be seen in Figure 1 (top) and Figure 2. Closest to ours is the concurrent work of [15]; however, as mentioned above, their approach is not fully automatic due to requiring a captioned source motion. Their keyframe selection heuristic further limits the applicability to certain edit types. Moreover, due to the unavailability of an open-source code, we do not provide comparisons in this paper.
@@
KIT-ML [39] 3911 1623 motion description
BABEL [41] 10881 1347 motion description, action
HumanML3D [17] 14616 5371 motion description
PoseFix [10] 6157 x 2 1068 pose editing
MotionFix (ours) 4771 x 2 1280 motion editing
Table 1: Comparison with existing datasets: MotionFix is the first dataset supporting the task of text-based motion editing.
3D human & language datasets. The progress in controlling 3D humans with text has been driven by new datasets that pair 3D humans with language descriptions. In Table 1, we summarize the three most popular motion description datasets. KIT [39] is the first source of such data containing textual annotations for $\sim$4k motion sequences. To obtain sufficient data to train deep networks, AMASS [32] unifies several MoCap collections in the SMPL body format [31]; however, it lacks language descriptions. This is addressed by BABEL [41] and HumanML3D [17], which concurrently collect semantic annotations in the form of action categories and/or textual descriptions. While KIT, BABEL, HumanML3D enable text-to-motion generation training, they do not support editing. On the other hand, as can be next seen in Table 1, PoseFix [10] provides pose editing triplets, but does not support motion editing. Our MotionFix dataset supports motion editing training, while being at a similar scale to PoseFix in terms of the number of triplets and the vocabulary of edit texts.
3 The New MotionFix Dataset
Figure 2: Dataset samples: We display source motions (red) overlaid with target motions (green) from our MotionFix dataset, together with their corresponding text annotations.
@@ 3 The New MotionFix Dataset
We make use of a motion embedding space to find motion pairs that are similar. Specifically, we employ a recent text-to-motion retrieval model TMR [37]. TMR is trained with a contrastive loss on the latent space of motions and texts, and reports state-of-the-art results for text-motion retrieval. We observe that such a model, by design, has the ability to produce latent motion representations that, for a given motion, ranks the semantically close ones nearby in the embedding space. We then use TMR perform motion-to-motion retrieval. This is in similar spirit to using CLIP [42] for image similarity. We construct our dataset by finding such motion pairs from the AMASS MoCap collection [32].
@@
Once we curate a list of candidate motion pairs, we give them to annotators from Amazon Mechanical Turk (AMT). In the annotation interface, along with the instructions, we give representative examples with multiple plausible edit texts for each motion pair. We allow the option to skip motion pairs if they are too similar (no difference to describe), or if they are too different (no easy way to describe the difference). Quantitatively, 7% and 55% of the pairs were considered too similar or too different, respectively. For the remaining pairs, we found that the majority are suitable candidates for editing.
@@
The resulting MotionFix dataset contains $4771$ triplets of source-target motions and text annotations, as shown in Table 1. In contrast to previous motion description datasets that provide text-motion pairs [39, 17, 41], our dataset enables training for motion editing, by also including a source motion. MotionFix is similar in spirit to PoseFix [10], but our labels describe the difference between dynamic motions, as opposed to static poses. Our dataset involves unrestricted edits, leading to different edit types such as spatial edits “throw from higher”, temporal subtraction of actions “start standing not bent down”, mixture of both “bend down a bit more, stand up faster”, and repetitions with adjustments of the whole body motion “do one more repetition and extend arms and legs wider apart”. We include several visual examples in Figure 2 that show body part editing (“keep arms at shoulder height”), directional (“move in the other direction”) and temporal (“start crawling earlier”) changes. Detailed statistics regarding the texts and motions can also be viewed in Section A of the appendix.
4 Text-Driven Motion Editing Diffusion Model

### v3

@@ Abstract.
The focus of this paper is 3D motion editing. Given a 3D human motion and a textual description of the desired modification, our goal is to generate an edited motion as described by the text. The challenges include the lack of training data and the design of a model that faithfully edits the source motion. In this paper, we address both these challenges. We build a methodology to semi-automatically collect a dataset of triplets in the form of (i) a source motion, (ii) a target motion, and (iii) an edit text, and create the new MotionFix dataset. Having access to such data allows us to train a conditional diffusion model, TMED, that takes both the source motion and the edit text as input. We further build various baselines trained only on text-motion pairs datasets, and show superior performance of our model trained on triplets. We introduce new retrieval-based metrics for motion editing, and establish a new benchmark on the evaluation set of MotionFix. Our results are encouraging, paving the way for further research on finegrained motion generation. Code, models and data are available at our project website.
Motion Editing, Motion from Instructions
journalyear: 2024copyright: rightsretainedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687559isbn: 979-8-4007-1131-2/24/12submissionid: 114ccs: Computing methodologies Motion processing
@@
Figure 1. Our text-driven motion diffusion model (TMED) enables 3D human motion editing from natural language descriptions. To train this model, we introduce a semi-automatically collected dataset MotionFix that contains diverse types of editing such as modifying body parts, changing certain moments of the motion, or editing the speed or the style.
[…]
Samples from the MotionFix dataset (top) and example motion editing results from our TMED model (bottom).
1. Introduction
@@
Human motion control is an essential component of the animation pipeline, and involves creating a motion, as well as editing it until the motion matches the desired outcome. Text descriptions have emerged as one of the prominent ways to control motion generation (Petrovich et al., 2022; Tevet et al., 2023; Guo et al., 2022a; Zhang et al., 2023a). However, due to the inherent ambiguity in high-level language instructions, the resulting generation may not necessarily correspond to the motion one has in mind. An animator may then need to edit the motion further. Motion editing is non-trivial, arguably more complex than static pose editing, and may involve multiple types of instructions, such as changing the speed of a motion, modifying the repetitions for cyclic actions, adjusting the posture of a particular body or modifying a certain temporal segment of a motion. In this work, we aim, given an initial source motion and an edit description, to generate a new motion that follows the source motion and edits it according to the text instruction.
@@
There are existing approaches that can modify body coordinates (Karunratanakul et al., 2023) and lower/upper limbs (Zhang et al., 2023a; Tevet et al., 2023). However, they require manual selection of the body parts, which prevents making edits beyond local modifications such as the speed of the overall action. On the other hand, some methods have been proposed to add more fine-grained control to text-to-motion generation; examples include temporal (Athanasiou et al., 2022; Zhang et al., 2023c) and spatial (Athanasiou et al., 2023; Zou et al., 2024; Goel et al., 2024) compositions. These methods can be repurposed for editing, but would be limited to addition or subtraction of actions. Consequently, all these works are limited to specific types of edits. Instead, our work considers unrestricted edits described by language instructions. To this end, we collect a manually annotated dataset, MotionFix, that supports training generative models for the task of text-driven motion editing.
@@
Constructing a dataset for 3D human motion editing is non-trivial. In contrast to text-based image editing, where methods such as InstructPix2Pix (Brooks et al., 2023) exploit large text-to-image generation models (Rombach et al., 2022) to automatically create training data, there exists no 3D motion generation model which generalizes faithfully to unrestricted text inputs. Besides, dynamic edits can be more complex than static image edits. In fact, PoseFix (Delmas et al., 2023) is a successful example of a 3D human body pose editing dataset; however, the difference between static poses are mapped to text in a rule-based manner using joint distances, and this would not be applicable to dynamic motions. In this work, we take a different route and mine existing motion capture (MoCap) datasets to find suitable motion pairs automatically, for which the differences are then manually described by typing text. By not relying on a generative model, we ensure motion quality; and by annotating the text (which is relatively fast), we achieve unrestricted edits.
@@
The key challenge in our semi-automatic data curation pipeline is how to find motion pairs that are similar enough for a meaningful and concise edit text to describe the difference. The difference between the source and target motions should not be too large to avoid the annotator typing a too-complex text or describing entirely the target motion, discarding the source. The motions should have similarities for the annotator to potentially make reference to the source. Our solution is to employ the recent TMR motion embedding space (Petrovich et al., 2023) that effectively captures semantics, as well as sufficient details for the body dynamics, thanks to its contrastively and generatively trained motion encoder. To form our candidate pairs for annotation, for each motion in a large MoCap collection (Mahmood et al., 2019), we retrieve the top-ranked motions according to their embedding similarity. Using crowdsourcing, we collect textual annotations for these pairs. The resulting dataset, MotionFix, is the first text-based motion editing dataset, which contains different types of edits as can be seen in Figure 1 (top). Some edits involve a specific body part (e.g., the hand should “make the circular motion a bit wider”), others alter the overall body dynamics (e.g., “make the circle bigger and walk faster” when walking in a circle).
@@
To benchmark our model and compare against baselines, we introduce new metrics on the evaluation set of MotionFix. Following the commonly adopted retrieval-based metrics in text-to-motion generation benchmarks (Guo et al., 2022a), we perform motion-to-motion retrieval and check how often the ground-truth target motion is in the top ranks. We also report the ranking of the source motion to evaluate the proximity to the source. While this metric should not be too high – otherwise there would be no edit – it gives an intuition on whether the generated motion deviates too much from the source. Our experiments demonstrate that our conditional model trained on triplets generates motions that are closer to the target, compared to strong baselines we build on top of state-of-the-art text-to-motion generation methods, which have only access to text-motion pairs for training.
@@ 2. Related Work
3D motion generation from text. In contrast to the relatively mature areas of text-to-image generation (Rombach et al., 2022) and text-based image editing (Brooks et al., 2023), language-based 3D human motion generation is at its infancy. Initial work employs VAEs (Kingma and Welling, 2014) with action label conditioning using a small set of categories (Guo et al., 2020; Petrovich et al., 2021). With the introduction of recent text-motion datasets (Punnakkal et al., 2021; Guo et al., 2022a; Lin et al., 2023), there has been increased interest in conditioning the generation on free-form language inputs (Petrovich et al., 2022; Guo et al., 2022a, b; Tevet et al., 2022; Uchida et al., 2024; Zou et al., 2024; Guo et al., 2024). Recently, diffusion models (Ho et al., 2020) have been successfully integrated (Zhang et al., 2023a; Tevet et al., 2023; Wan et al., 2023; Shafir et al., 2024; Xie et al., 2024), producing state-of-the-art results in text-to-motion generation.
@@
Several works focus on increased controllability in motion generation, going beyond a single textual input. Examples include enabling temporal compositionality (a series of motions) (Athanasiou et al., 2022; Lee et al., 2022; Shafir et al., 2024), spatial compositionality (Athanasiou et al., 2023; Zhang et al., 2023a) (simultaneous motions), and a unified framework of timeline control (Petrovich et al., 2024). Diffusion-based models were shown to be suitable for fine-grained local control such as joint trajectories (Karunratanakul et al., 2023; Xie et al., 2024; Shafir et al., 2024) or keyframes (Cohan et al., 2024). Our work is in similar spirit in terms of providing more control to users; however, in contrast to the above, our focus is motion editing using language.
@@
Language-based human body editing. There are numerous traditional methods for editing 3D humans to generate movies (Catmull, 1972), imposing space or time constraints (Cohen, 1992), interactively generating motions using procedural animation (Lee and Shin, 1999; Perlin, 1995) or physics-based approaches (Popović and Witkin, 1999). Recent work using language can be grouped into pose (Kim et al., 2021; Delmas et al., 2023) or motion (Fieraru et al., 2021; Goel et al., 2024) editing. In FixMyPose (Kim et al., 2021), the focus is on editing athletic human poses in synthetic images. In (Delmas et al., 2023), a text-based 3D human pose editing method is developed, enabled through the collection of the PoseFix dataset containing language descriptions of differences between pairs of poses. PoseFix builds on the previous work of PoseScript (Delmas et al., 2022), where a dataset of pose descriptions are automatically collected through a rule-based approach. Unlike PoseFix that concentrates on static poses, our MotionFix dataset involves dynamic motions where the space of possible edits are much larger, necessitating a different approach to data collection.
@@
In terms of dynamic bodies, current motion editing approaches can be separated into three categories: (a) Style-based editing or motion style transfer exploits datasets that contain a small set of style labels such as ‘angry’ and ‘old’ (Aberman et al., 2020; Mason et al., 2022; Kobayashi et al., 2023). This line of work, focuses mostly on copying the style of one motion onto another, typically performing the same action. (b) Part-based editing considers selecting a subset of the body. MDM (Tevet et al., 2023) shows the potential of diffusion models to edit the upper/lower body by text-conditioned motion inpainting. Similarly, MotionDiffuse (Zhang et al., 2023a) and FLAME (Kim et al., 2023) manually specify body parts to edit them with text. More recently, CoMo (Huang et al., 2024) and FineMoGen (Zhang et al., 2023b) use LLMs to produce edit texts and demonstrate promising results for part editing. (c) Among heuristic-based approaches (Fieraru et al., 2021; Goel et al., 2024), AIFit (Fieraru et al., 2021) can edit domain-specific exercise poses from a pre-defined grammar and is focused on a limited set of cyclic motions from their Fit3D dataset. Iterative Motion Editing (Goel et al., 2024) relies on captioned source motions, which are passed through an LLM along with a pre-defined set of ‘Motion Editing Operators’ (MEOs), to detect which joints and frames should be edited. A pre-trained diffusion model is then used to infill these locations. In contrast with the prior work, we do not focus on a specific type of motion editing or any heuristics. Our MotionFix contains diverse edits as can be seen in Figure 1 (top) and Figure 2. Closest to ours is the concurrent work of (Goel et al., 2024); however, as mentioned above, their approach is not fully automatic due to requiring a captioned source motion. Their keyframe selection heuristic further limits the applicability to certain edit types. Moreover, due to the unavailability of an open-source code at the time of writing this paper, we do not provide comparisons in this work.
@@
KIT-ML (Plappert et al., 2016) 3911 1623 motion description
BABEL (Punnakkal et al., 2021) 10881 1347 motion description, action
HumanML3D (Guo et al., 2022a) 14616 5371 motion description
PoseFix (Delmas et al., 2023) 6157 x 2 1068 pose editing
MotionFix (ours) 6730 x 2 1479 motion editing
Table 1. Comparison with existing datasets: MotionFix is the first dataset supporting the task of text-based motion editing.
3D human & language datasets. The progress in controlling 3D humans with text has been driven by new datasets that pair 3D humans with language descriptions. In Table 1, we summarize the three most popular motion description datasets. KIT (Plappert et al., 2016) is the first source of such data containing textual annotations for $\sim$4k motion sequences. To obtain sufficient data to train deep networks, AMASS (Mahmood et al., 2019) unifies several MoCap collections in the SMPL body format (Loper et al., 2015); however, it lacks language descriptions. This is addressed by BABEL (Punnakkal et al., 2021) and HumanML3D (Guo et al., 2022a), which concurrently collect semantic annotations in the form of action categories and/or textual descriptions. While KIT, BABEL, HumanML3D enable text-to-motion generation training, they do not support editing. On the other hand, as can be next seen in Table 1, PoseFix (Delmas et al., 2023) provides pose editing triplets, but does not support motion editing. Our MotionFix dataset supports motion editing training, while being at a similar scale to PoseFix in terms of the number of triplets and the vocabulary of edit texts.
3. The New MotionFix Dataset
Figure 2. Dataset samples: We display source motions (red) overlaid with target motions (green) from our MotionFix dataset, together with their corresponding text annotations.
[…]
Dataset Samples
@@ 3. The New MotionFix Dataset
We make use of a motion embedding space to find motion pairs that are similar. Specifically, we employ a recent text-to-motion retrieval model TMR (Petrovich et al., 2023). TMR is trained with a contrastive loss on the latent space of motions and texts, and reports state-of-the-art results for text-motion retrieval. We observe that such a model, by design, has the ability to produce latent motion representations that, for a given motion, ranks the semantically close ones nearby in the embedding space. We then use TMR to perform motion-to-motion retrieval. This is in similar spirit to using CLIP (Radford et al., 2021) for image similarity. We construct our dataset by finding such motion pairs from the AMASS MoCap collection (Mahmood et al., 2019).
@@
Once we curate a list of candidate motion pairs, we give them to annotators from Amazon Mechanical Turk (AMT). In the annotation interface, along with the instructions, we give representative examples with multiple plausible edit texts for each motion pair. We allow the option to skip motion pairs if they are too similar (no difference to describe), or if they are too different (no easy way to describe the difference). Quantitatively, 7% and 55% of the pairs were considered too similar or too different, respectively. For the remaining pairs, we found that the majority are suitable candidates for editing. Our annotation interface can be visualized in the supplementary materials (sup. mat.).
@@
The resulting MotionFix dataset contains $6730$ triplets of source-target motions and text annotations. We partition the data into train/validation/test splits randomly with 80%/5%/15% ratios, and obtain 5387/330/1013 triplets for each split, respectively. As shown in Table 1, in contrast to previous motion description datasets that provide text-motion pairs (Plappert et al., 2016; Guo et al., 2022a; Punnakkal et al., 2021), our dataset enables training for motion editing, by also including a source motion. MotionFix is similar in spirit to PoseFix (Delmas et al., 2023), but our labels describe the difference between dynamic motions, as opposed to static poses. Our dataset involves unrestricted edits, leading to different edit types such as spatial edits “throw from higher”, temporal subtraction of actions “start standing not bent down”, mixture of both “bend down a bit more, stand up faster”, and repetitions with adjustments of the whole body motion “do one more repetition and extend arms and legs wider apart”. We include several visual examples in Figure 2 that show body part editing (“keep arms at shoulder height”), directional (“move in the other direction”) and temporal (“start crawling earlier”) changes. We provide dynamic video examples in our webpage [^1] https://motionfix.is.tue.mpg.de through the supplementary video and the data exploration interface. Detailed statistics regarding the texts and motions can also be viewed in Section A of the appendix.
4. Text-Driven Motion Editing Diffusion Model